Break-Even Sales

Currently, the unit selling price of a product is $370, the unit variable cost is $300, and the total fixed costs are $1,078,000. A proposal is being evaluated to increase the unit selling price to $410.

a. Compute the current break-even sales (units).
units

b. Compute the anticipated break-even sales (units), assuming that the unit selling price is increased and all costs remain constant.
units
Answer #1

Requirement a

current break-even sales (units).= 15,400 Units

.Working

A Sales price per unit $                  370.00
B Variable cost per unit $                  300.00
C=A-B Contribution margin per unit $                    70.00
D Fixed cost $      1,078,000.00
E=D/C Break Even point in Unit Sales                15,400.00

.

Requirement 2

anticipated break-even sales (units) = 9800 Units

.Working

A Sales price per unit $                  410.00
B Variable cost per unit $                  300.00
C=A-B Contribution margin per unit $                  110.00
D Fixed cost $      1,078,000.00
E=D/C Break Even point in Unit Sales                   9,800.00